OpenStudy (bookworm14):

A teacher asks each student in the class to randomly choose a number from 1 to 50 and write it down. Jean and Serita each write down a number. Find the probability that both student’s choices will be greater than 35. Express your answer as a decimal rounded to the nearest thousandth.

OpenStudy (dan815):

there are 15 numbers greater than 35 that are 50 or less

OpenStudy (bookworm14):

yes, i got 0.3 but im not sure if that is the final answer considering there are 2 people drawing

OpenStudy (dan815):

15/50 is the chance jean will get over 35 15/50 is the chance serita will get over 35 (15/50)^2 is the chance they both get over 35

OpenStudy (dan815):

(15/50)^2=(3/10)^2=9/100

OpenStudy (dan815):

9% or 0.09

OpenStudy (bookworm14):

sorry i my mind wondered off in another one of my math questions, and i understand how you got that thank you!